# Mass hierarchies in gauge theory with two index symmetric representation matter

### v1

@@ II The Holographic Model
Note the only interplay between the two representations is through their contributions to $\Delta m^{2}_{R}$. The five-dimensional coupling may be obtained by matching to the UV vector-vector correlator Erlich:2005qh , and is given by
@@
The vector meson decay constant is then given by the overlap term between the meson and the external source
@@ III.1 $N_{c}=5$ Theory
In Figure 2 we show the resulting field $L_{15}(\rho)$ in the holographic model (also in blue) - it can be thought of as the effective fermion mass as a function of RG scale. It shows chiral symmetry breaking, bending away from $L=0$ in the infra-red. As hinted at above the IR value of $L_{15}(\rho)$ is much smaller than the scale where the $\gamma=1$/BF bound is violated.
@@
We can now solve for the $L(\rho)$ function for the 5 representation. We display this in green in Figure 2. We can see there is a gap of a factor of 10.7 between the IR mass values for the two representations. This is of order but not quite as large as the factor of 15 we naïvely expected. Here there is not a large gap between the BF bound violation scale and the emergent IR mass because the running is much faster once the 15 have decoupled (although we have chosen $N_{f}^{F}$ to slow this running to generate the gap). The gap size is therefore in line with the expectations of the running analysis of Evans:2020ztq although both scales are significantly lower than expected from the $\gamma_{15}=1$ criteria.
@@
The $\sigma$ meson in the symmetric representation sector (green) is the lightest state reflecting the walking in the high energy theory. Its decay constant initially grows, before falling off at high $N_{f}^{F}$ in the more walking theory.

### v2

@@ II The Holographic Model
Note the only interplay between the two representations is through their contributions to $\Delta m^{2}_{R}$. The five-dimensional coupling may be obtained by matching to the UV vector-vector correlator Erlich:2005qh ; DaRold:2005mxj , and is given by
@@
Note these gauge theory results are all gauge invariant and scheme independent as discussed in Appelquist:1996dq ; Ryttov:2010iz in the context of the conformal window.
@@
The vector meson decay constant is then given by the
overlap term between the meson and the external source
@@ III.1 $N_{c}=5$ Theory
In Figure 2 we show the resulting field $L_{15}(\rho)$ in the holographic model (also in blue) - it can be thought of as the
@@
effective fermion mass as a function of RG scale. It shows chiral symmetry breaking, bending away from $L=0$ in the infra-red. As hinted at above the IR value of $L_{15}(\rho)$ is much smaller than the scale where the $\gamma=1$/BF bound is violated.
@@
We can now solve for the $L(\rho)$ function for the 5 representation. We display this in green in Figure 2. We can see there is a gap of a factor of 10.7 between the IR mass values for the two representations. This is of order but not quite as large as the factor of 15 we naïvely expected. Here there is not a large gap between the BF bound violation scale and the emergent IR mass because
@@
the running is much faster once the 15 have decoupled (although we have chosen $N_{f}^{F}$ to slow this running to generate the gap). The gap size is therefore in line with the expectations of the running analysis of Evans:2020ztq although both scales are significantly lower than expected from the $\gamma_{15}=1$ criteria.
@@
The $\sigma$ meson in the symmetric representation sector (green) is the lightest state reflecting the walking in the high energy theory. Its decay constant initially grows,
@@
before falling off at high $N_{f}^{F}$ in the more walking theory.